# Homeomorphism groups of basilica, rabbit and airplane Julia sets

### v1

@@ 2.1 Topological characterization of the airplane
In the first case, a connected component $X$ of $\mathcal{A}\setminus\{p\}$ includes both subarcs $[x,p)$ and $(p,y]$ of $A$. In particular it contains $x$ and $y$ and is path-connected (because it is a Peano continuum), so $X$ includes an arc $A^{*}$ between $x$ and $y$. Thus, the subset $A^{*}\cup[x,p)\cup[y,p)\cup\{p\}$ of $X\cup\{p\}$ includes a circle that meets $C$ in $p$, which contradicts condition 3 of B.
In the second case, a connected component $X$ of $\mathcal{A}\setminus\{p\}$ includes both $C\setminus\{p\}$ and a subarc $[x,p)$ or $(p,y]$ of $A$. If we suppose for example that $[x,p)\subseteq X$, then $X$ includes an arc $[x,z]$ between $x$ and some $z\in C\setminus\{p\}$. Let $[z,p]$ be any arc between $z$ and $p$ that is included in the circle $C$. The subset $[x,z]\cup[z,p]\cup(p,x]$ includes a circle in $X\cup\{p\}$ that meets $C$ in $p$ and is distinct from $C$, which contradicts condition 3 of B. ∎
@@
$C(x,y)=\{C\text{ circle }\mid x\in C\text{ or }y\in C\}\cup\{C\text{ circle }\mid C\text{ separates }x\text{ and }y\}.$
@@
Given an airplane $\mathcal{A}$, the collection of all connected components of $\mathcal{A}$ minus finitely many of its circles and circular cut points is a subbasis for the topology of $\mathcal{A}$.
@@ 2.2 The dendrite of circles of an airplane
Given a subset $S\subseteq\mathbf{N}_{\geq 3}\cup\{\infty\}$, the Ważewski dendrite associated to $S$ is a dendrite such the order of any branch point belongs to $S$ and that, for each $s\in S$, any arc of the dendrite contains points of order $s$. Such dendrite is usually denoted by $D_{S}$ or simply by $D_{n}$ when $S=\{n\}\subseteq\mathbf{N}_{\geq 3}\cup\{\infty\}$. By [CD94, Theorem 6.2], for each distinct $S$ there is a unique Ważewski dendrite $D_{S}$ up to homeomorphism, so we can refer to $D_{S}$ as the Ważewski dendrite of order $S$. Mimicking the construction of the tree of circles for rabbits, we now associate a copy of $D_{\infty}$ to the airplane (this construction was previously outlined in [Tar23, Section 7.3]).
@@ 2.4 Kaleidoscopic groups
$\sigma_{g}(i)\coloneqq k_{g(b)}\circ g\circ k_{b}^{-1}.$
@@ 2.5 Uniqueness of the airplane
[…]
@@
For each circle $C$ of an airplane, each cut point $x\in C$ corresponds to a unique wing $U_{x}$ not including $C$. For $C\subseteq\mathcal{A}_{1}$, the image $\pi_{1}(U_{x})$ is a branch of $Q_{1}$ around the image $b_{1}=\pi_{1}(C)$. The homeomorphism $\varphi$ maps such a branch of $Q_{1}$ to a branch of $Q_{2}$ around $b_{2}=\varphi(b_{1})$ which is the image of a unique wing $U_{y}$ associated to a circular cut point on the circle $\pi_{2}^{-1}(\varphi\circ\pi_{1}(C))$, so we can extend $\psi$ to the set of all cut points that lie on circles. Since $\varphi$ induces a bijection that preserves the separation relation between branches around a branch point $b$ to branches around the branch point $\varphi(b)$, the map $\psi$ restricted to cut points on a given circle $C$ also preserves the separation relation and thus extends uniquely to a continuous map between $C$ and the circle that contains the images of cut points in $C$. Doing that for all circles $C\subset\mathcal{A}_{1}$ extends $\psi$ to a bijection between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $\psi\circ\pi_{1}=\pi_{2}\circ\psi$. It now remains to show that $\psi$ is a homeomorphism.
Let $x\in\mathcal{A}_{1}$ be a point that does not lie on a circle. Then $\psi(x)$ does not lie on a circle. By conditions 1 and 2 of B, the collection of wings that contain $\psi(x)$ satisfies the hypothesis of Proposition 1.9 and is thus a basis of neighborhoods at $\psi(x)$. For any wing $W$ that contains $\psi(x)$, its preimage $\psi^{-1}(W)$ is a wing that contains $x$, which is open. Thus, $\psi$ is continuous at points $x$ that do not lie on circles.
@@ 2.6 Homeomorphism group of the airplane
The embedding $\Pi$ of Proposition 2.13 is a surjective.
@@ 3.1 Laminations of quadratic Julia sets
In Appendix B, we prove that $n$-rabbits and the airplane are homeomorphic to the Julia set of some quadratic polynomial $f_{c}(z)=z^{2}+c$, where $c\in\mathbf{C}$ depends on the space considered. Identifying our spaces with a Julia set $J_{c}$, we obtain a continuous surjective map $\varphi\colon\mathbf{S}^{1}\to J_{c}$ called the Carathéodory loop of $J_{c}$ (see for example [DH84a, Exposé II]). Let us fix one of the two cyclic orders on $\mathbf{S}^{1}$.
@@
Let us recall that a point $c\in\mathbf{C}$ belongs to the Mandelbrot set $\mathcal{M}$ if the Julia set of $f_{c}$ is connected, which is equivalent to the fact that the orbit of the origin is bounded. The polynomial $f_{c}$ is (sub)hyperbolic if it is (sub)expanding on its Julia set $J_{c}$. The filled-in Julia set $K_{c}$ is the set of points $z\in\mathbf{C}$ with bounded orbits. The boundary $\partial K_{c}$ is $J_{c}$ and the connected components of $\overset{\circ}{K_{c}}$ are topological disks which are in bijection with the circles of $J_{c}$ (their boundaries). When $f_{c}$ is hyperbolic, both $K_{c}$ and $J_{c}$ are connected and locally connected. See [DH84a, Exposé III] for more details.
@@ 3.1.1 Orientations and projections on circles and cut points
Similarly, for each circle $C$ of $J_{c}$, the orientation $O_{\mathbf{S}^{1}}$ of $\mathbf{S}^{1}$ induces a cyclic order $O_{\varphi}^{C}$ on $C$: if $\varphi(x)$, $\varphi(y)$ and $\varphi(z)$ belong to $C$, we set $O_{\varphi}^{C}(\varphi(x),\varphi(y),\varphi(z))$ as $+$ or -- depending on the orientation of the points $x$, $y$ and $z$ on $\mathbf{S}^{1}$. Note that each cyclic order $O_{\varphi}^{C}$ on a circle $C$ of $J_{c}$ corresponds to one of the two cyclic orders that define the separation relation $S$ on $C$ which we used in Sections 1.6, LABEL: and 2.6 to determine the local actions $\mathrm{Aut}(S)$ of the homeomorphism groups of the rabbits and the airplane.
For each circle $C$ of $J_{c}$ and each point $p$ of $J_{c}$ there is a unique point that is the projection of $p$ onto $C$ [DH84a, Exposé II, Proposition 4]. We denote it by $P_{C}(p)$.
@@
Let $E_{x}$, $E_{y}$ and $E_{z}$ be the external rays of $K_{c}$ with arguments $x$, $y$ and $z$, respectively. Consider legal arcs $[\varphi(x),P_{C}(\varphi(x))]$, $[\varphi(y),P_{C}(\varphi(y))]$ and $[\varphi(z),P_{C}(\varphi(z))]$ in $K_{c}$. Since $P_{C}(\varphi(x))$, $P_{C}(\varphi(y))$ and $P_{C}(\varphi(z))$ are distinct points of $C$, these arcs lie in distinct components of $K_{C}\setminus D$, where $D$ is the disk in $K_{c}$ such that $C=\partial D\subseteq J_{c}$. So $E_{x}\cup[\varphi(x),P_{C}(\varphi(x))]$, $E_{y}\cup[\varphi(y),P_{C}(\varphi(y))]$ and $E_{z}\cup[\varphi(z),P_{C}(\varphi(z))]$ are pairwise disjoint and the statement about the cyclic order follows from the definition of $O_{\varphi}^{C}$. ∎
@@ 3.2 Identification of the lamination automorphism groups
Given $x$, $y$ and $z$ distinct vertices of a tree or points of a dendrite, their center is the unique point in the intersection $[x,y]\cap[y,z]\cap[z,x]$, where $[p,q]$ denotes the unique path in the tree or arc in the dendrite that joins $p$ and $q$.
@@
Ultimately, the action of $\mathrm{Homeo}^{+}(J)$ preserves the cyclic order on the dense subset $R$ of $\mathbf{S}^{1}$. By [Duc23a, Lemma 3.5], it thus extends to a unique element of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$. By construction, it preserves the laminational equivalence relation $\mathcal{L}J$, so we have an embedding of $\mathrm{Homeo}^{+}(J)$ into $\mathrm{Aut}(\mathcal{L}J)$. Conversely, any element of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ that preserves $\mathcal{L}J$ induces a homeomorphism of the quotient space $\mathbf{S}^{1}/\mathcal{L}J=J$. By construction of the cyclic orders $O_{\varphi}^{p}$ and $O_{\varphi}^{C}$, such homeomorphism has local actions in $\mathrm{Cyc}(n)$ or $\mathrm{Aut}(O)$. The fact that this correspondence is an isomorphism of topological groups follows from the fact that the Polish topologies coincide with the pointwise convergence topologies on $R$ and $\varphi(R)$. ∎
@@
It is not hard to see that $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ (respectively $\mathrm{Homeo}^{+}(\mathcal{A})$) has infinite index and is not normal in $\mathrm{Homeo}(\mathcal{R}_{n})$ (respectively $\mathrm{Homeo}(\mathcal{A})$. This contrasts with the fact that $\mathrm{Aut}(O)$ is an index-$2$ normal subgroup of $\mathrm{Aut}(S)$ (Lemma A.8).
@@ 3.3 Simplicity of commutator subgroups
Figure 9: The $3$-coloring of the circles of $\mathcal{R}_{3}$.
@@
The basilica and airplane rearrangement groups $T_{B}$ and $T_{A}$ from [BF15, BF19, Tar24b] are countable dense subgroups of the groups $\mathrm{Homeo}^{+}(\mathcal{B})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$, respectively. For the basilica this was shown in [Ner24] and for the airplane a proof is sketched in [Tar23, Subsection 7.3].
Note that the behavior of $\mathrm{Homeo}^{+}(\mathcal{B})$ described in Theorem 3.9 is shared by $T_{B}$, since its commutator subgroup $[T_{B},T_{B}]$ is a simple index-$2$ normal subgroup consisting of those elements of $T_{B}$ that preserve the $2$-coloring depicted in Figure 8 [BF15, Section 8]. Instead, unlike $\mathrm{Homeo}^{+}(\mathcal{A})$, the airplane rearrangement group $T_{A}$ has infinite-index simple commutator subgroup [Tar24b, Corollary 7.4 and Thoerem 7.8].
@@
    A rearrangement group of the basilica that does not preserve the orientation of the circles can be built by switching orientation of the edge attached to the initial vertex of the replacement graph of the basilica replacement system. The same arguments employed in [Ner24, Tar23] show that it is dense in $\mathrm{Homeo}(\mathcal{B})$. Does it have a simple index-$2$ commutator subgroup?
@@ 4.1 Patchwork
Since most of the statements are essentially the same in both settings, we will let $T$ be a biregular tree and $D$ be a Ważewski dendrite. Given $p$ and $q$ vertices of a tree or points of a dendrite, we will denote by $[p,q]$ either the unique path in the tree (as a set of vertices) or the unique arc in the dendrite (as a subset of the dendrite) between $p$ and $q$. Recall that, given $x$, $y$ and $z$ distinct vertices of $T$ or points of $D$, their center is the unique point in the intersection $[x,y]\cap[y,z]\cap[z,x]$.
@@
  • Given subsets $F$ and $F^{\prime}$ of $V(T)$, a partial tree homomorphism $F\to F^{\prime}$ is a bijection that is a restriction of a tree isomorphism $[F]\to[F^{\prime}]$.
@@
It is easy to see that any partial tree homomorphism $F\to F^{\prime}$ is the restriction of a unique tree isomorphism $[F]\to[F^{\prime}]$.
@@
If $F$ is a center-closed subset of a dendrite $D$, then each element of $\Omega_{F}$ has either one or two points of $F$ on its topological boundary. Indeed, if $\partial A\supseteq\{x,y,z\}$ for some $A\in\Omega_{F}$, then the center of $x$, $y$ and $z$ would belong to $F$ and so $A$ could not be a connected component of $D\setminus F$.
@@
Suppose that $F$ and $F^{\prime}$ are subsets of $V(T)$ or $D$ (center-closed in the case of $D$) and let $f\colon F\to F^{\prime}$ be a partial homomorphism. Assume the following:
@@
Define a map $h\colon V(T)\to V(T)$ or $D\to D$ by setting $h(x)=h_{A}(x)$ when $x\in A\cup$ for some $A\in\Omega_{F}$ and $h(x)=f(x)$ for all $x\in F$. Then $h$ is an automorphism of $T$ or a homeomorphism of $D$ that agrees with $f$ on $F$.
@@
For dendrites, this is Lemma 4.3 from [DMW19] when $F$ and $F^{\prime}$ are finite. When they are infinite, the continuity of $h$ is proved by observing that $h$ maps each arc to an arc and applying [DMW19, Proposition 2.4].
Let us now consider the case of a tree. In this case, first note that $h$ is a well-defined bijection because $\Omega_{F}\cup\{F\}$ is a partition of $V(T)$. Then we only need to check that the bijection $h$ of the vertices of $T$ is truly a tree automorphism, i.e., that $\{v,w\}\in E(T)$ if and only if $\{h(v),h(w)\}\in E(T)$. When both $v$ and $w$ belong to $F$, this follows from the fact that $h|_{F}=f|_{F}$ and that $f$ descends from an isomorphism $[F]\to[F^{\prime}]$. On the other hand, when either or both vertices belong to the complement of $F$, since they are at distance $1$ they must belong to a unique common $A\cup\partial A$, so the conclusion follows from the fact that each $h_{A}$ is a tree automorphism defined on $A\cup F$. ∎

### v2

@@ 2.1 Topological characterization of the airplane
In the first case, a connected component $X$ of $\mathcal{A}\setminus\{p\}$ includes both subarcs $[x,p)$ and $(p,y]$ of $A$. In particular it contains $x$ and $y$ and is path-connected (because it is a connected and open in $\mathcal{A}$), so $X$ includes an arc $A^{*}$ between $x$ and $y$. Thus, the subset $A^{*}\cup[x,p)\cup[y,p)\cup\{p\}$ of $X\cup\{p\}$ includes a circle that meets $C$ in $p$, which contradicts condition 3 of B.
In the second case, a connected component $X$ of $\mathcal{A}\setminus\{p\}$ includes both $C\setminus\{p\}$ and a subarc $[x,p)$ or $(p,y]$ of $A$. If we suppose for example that $x,p)\subseteq X$, then $X$ includes an arc $[x,z]$ between $x$ and some $z\in C\setminus\{p\}$. Let $[z,p]$ be any arc between $z$ and $p$ that is included in the circle $C$. The subset $x,z]\cup[z,p]\cup(p,x]$ includes a circle in $X\cup\{p\}$ that meets $C$ in $p$ and is distinct from $C$, which contradicts condition 3 of B. ∎
@@
[…]
@@
Given an airplane $\mathcal{A}$, the collection of all connected components of $\mathcal{A}$ minus finitely many of its circles and cut point that lie on circles is a subbasis for the topology of $\mathcal{A}$.
@@ 2.2 The dendrite of circles of an airplane
Given a subset $S\subseteq\mathbf{N}_{\geq 3}\cup\{\infty\}$, the Ważewski dendrite associated to $S$ is a dendrite such the order of any branch point belongs to $S$ and that, for each $s\in S$, any arc of the dendrite contains points of order $s$. Such dendrite is usually denoted by $D_{S}$ or simply by $D_{n}$ when $S=\{n\}\subseteq\mathbf{N}_{\geq 3}\cup\{\infty\}$. By [CD94, Theorem 6.2], for each distinct $S$ there is a unique Ważewski dendrite $D_{S}$ up to homeomorphism, so we can refer to $D_{S}$ as the Ważewski dendrite of order $S$. Mimicking the construction of the tree of circles for rabbits, we now associate a copy of $D_{\infty}$ to the airplane (this construction was previously outlined in [Tar25, Section 7.3]).
@@ 2.4 Kaleidoscopic groups
$\sigma_{g}(b)\coloneqq k_{g(b)}\circ g\circ k_{b}^{-1}.$
@@ 2.5 Uniqueness of the airplane
$\mathcal{A}_{1}$$\mathcal{A}_{2}$$D_{\infty}$$D_{\infty}$$\scriptstyle{\psi}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\varphi}$
@@
For each circle $C$ of an airplane, each cut point $x\in C$ corresponds to a unique wing $U_{x}$ not including $C$. For $C\subseteq\mathcal{A}_{1}$, the image $\pi_{1}(U_{x})$ is a branch of $Q_{1}$ around the image $b_{1}=\pi_{1}(C)$. The homeomorphism $\varphi$ maps such a branch of $Q_{1}$ to a branch of $Q_{2}$ around $b_{2}=\varphi(b_{1})$ which is the image of a unique wing $U_{y}$ associated to a circular cut point on the circle $\pi_{2}^{-1}(\varphi\circ\pi_{1}(C))$, so we can extend $\psi$ to the set of all cut points that lie on circles. Since $\varphi$ induces a bijection that preserves the separation relation between branches around a branch point $b$ to branches around the branch point $\varphi(b)$, the map $\psi$ restricted to cut points on a given circle $C$ also preserves the separation relation and thus extends uniquely to a continuous map between $C$ and the circle that contains the images of cut points in $C$. Doing that for all circles $C\subset\mathcal{A}_{1}$ extends $\psi$ to a bijection between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $\varphi\circ\pi_{1}=\pi_{2}\circ\psi$. It now remains to show that $\psi$ is a homeomorphism.
Let $x\in\mathcal{A}_{1}$ be a point that does not lie on a circle. Then $\psi(x)$ does not lie on a circle. By conditions 1 and 2 of B, the collection of wings that contain $\psi(x)$ satisfies the hypothesis of Proposition 1.9 and is thus a subbasis of neighborhoods at $\psi(x)$. For any wing $W$ that contains $\psi(x)$, its preimage $\psi^{-1}(W)$ is a wing that contains $x$, which is open. Thus, $\psi$ is continuous at points $x$ that do not lie on circles.
@@ 2.6 Homeomorphism group of the airplane
The embedding $\Pi$ of Proposition 2.13 is surjective.
@@ 3.1 Laminations of quadratic Julia sets
In Appendix B, we prove that $n$-rabbits and the airplane are homeomorphic to the Julia set of some quadratic polynomial $f_{c}(z)=z^{2}+c$, where $c\in\mathbf{C}$ depends on the space considered. Identifying our spaces with a Julia set $J_{c}$, we obtain a continuous surjective map $\varphi\colon\mathbf{S}^{1}\to J_{c}$ called the Carathéodory loop of $J_{c}$ (see for example [DH84, Exposé II]). Let us fix one of the two cyclic orders on $\mathbf{S}^{1}$.
@@
Let us recall that a point $c\in\mathbf{C}$ belongs to the Mandelbrot set if the Julia set of $f_{c}$ is connected, which is equivalent to the fact that the orbit of the origin is bounded. The polynomial $f_{c}$ is (sub)hyperbolic if it is (sub)expanding on its Julia set $J_{c}$. The filled-in Julia set $K_{c}$ is the set of points $z\in\mathbf{C}$ with bounded orbits. The boundary $\partial K_{c}$ is $J_{c}$ and the connected components of $\overset{\circ}{K_{c}}$ are topological disks which are in bijection with the circles of $J_{c}$ (their boundaries). When $f_{c}$ is hyperbolic, both $K_{c}$ and $J_{c}$ are connected and locally connected. See [DH84, Exposé III] for more details.
@@ 3.1.1 Orientations and projections on circles and cut points
Similarly, for each circle $C$ of $J_{c}$, the orientation $O_{\mathbf{S}^{1}}$ of $\mathbf{S}^{1}$ induces a cyclic order $O_{\varphi}^{C}$ on $C$: if $\varphi(x)$, $\varphi(y)$ and $\varphi(z)$ belong to $C$, we set $O_{\varphi}^{C}(\varphi(x),\varphi(y),\varphi(z))$ as ++ or - depending on the orientation of the points $x$, $y$ and $z$ on $\mathbf{S}^{1}$. Note that each cyclic order $O_{\varphi}^{C}$ on a circle $C$ of $J_{c}$ corresponds to one of the two cyclic orders that define the separation relation $S$ on $C$ which we used in Sections 1.6, LABEL: and 2.6 to determine the local actions $\mathrm{Aut}(S)$ of the homeomorphism groups of the rabbits and the airplane.
For each circle $C$ of $J_{c}$ and each point $p$ of $J_{c}$ there is a unique point that is the projection of $p$ onto $C$ [DH84, Exposé II, Proposition 4]. We denote it by $P_{C}(p)$. In the next lemma we use the classical notions of external rays and legal arcs; see [DH84a] for definitions (page 10 for external rays and page 19 for legal arcs, which are called allowable arcs in the reference).
@@
Let $E_{x}$, $E_{y}$ and $E_{z}$ be the external rays of $K_{c}$ with arguments $x$, $y$ and $z$, respectively. Consider legal arcs $\varphi(x),P_{C}(\varphi(x))]$, $\varphi(y),P_{C}(\varphi(y))]$ and $\varphi(z),P_{C}(\varphi(z))]$ in $K_{c}$. Since $P_{C}(\varphi(x))$, $P_{C}(\varphi(y))$ and $P_{C}(\varphi(z))$ are distinct points of $C$, these arcs lie in distinct components of $K_{C}\setminus D$, where $D$ is the disk in $K_{c}$ such that $C=\partial D\subseteq J_{c}$. So $E_{x}\cup[\varphi(x),P_{C}(\varphi(x))]$, $E_{y}\cup[\varphi(y),P_{C}(\varphi(y))]$ and $E_{z}\cup[\varphi(z),P_{C}(\varphi(z))]$ are pairwise disjoint and the statement about the cyclic order follows from the definition of $O_{\varphi}^{C}$. ∎
@@ 3.2 Identification of the lamination automorphism groups
Given $x$, $y$ and $z$ distinct vertices of a tree or points of a dendrite, their center is the unique point in the intersection $x,y]\cap[y,z]\cap[z,x]$, where $[p,q]$ denotes the unique path in the tree or arc in the dendrite that joins $p$ and $q$.
@@
Ultimately, the action of $\mathrm{Homeo}^{+}(J)$ preserves the cyclic order on the dense subset $R$ of $\mathbf{S}^{1}$. By [Duc23, Lemma 3.5], it thus extends to a unique element of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$. By construction, it preserves the laminational equivalence relation $\mathcal{L}J$, so we have an embedding of $\mathrm{Homeo}^{+}(J)$ into $\mathrm{Aut}(\mathcal{L}J)$. Conversely, any element of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ that preserves $\mathcal{L}J$ induces a homeomorphism of the quotient space $\mathbf{S}^{1}/\mathcal{L}J=J$. By construction of the cyclic orders $O_{\varphi}^{p}$ and $O_{\varphi}^{C}$, such homeomorphism has local actions in $\mathrm{Cyc}(n)$ or $\mathrm{Aut}(O)$. The fact that this correspondence is an isomorphism of topological groups follows from the fact that the Polish topologies coincide with the pointwise convergence topologies on $R$ and $\varphi(R)$. Indeed, the compact-open topology is always stronger. For the converse, it suffices to show it on all of $\mathbf{S}^{1}$. Here, since $R$ is dense in $\mathbf{S}^{1}$, one can one can cover $\mathbf{S}^{1}$ with arbitrarily small intervals whose end points lie in $R$. ∎
@@
By Lemma A.8 the group $\mathrm{Aut}(O)$ is an index-$2$ normal subgroup of $\mathrm{Aut}(S)$. This contrasts with the fact that $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ (respectively $\mathrm{Homeo}^{+}(\mathcal{A})$) has infinite index: note that, for any infinite path (arc) of the tree (dendrite) of circles, for each circle lying on it, one can find an element of $\mathrm{Homeo}(\mathcal{R}_{n})$ ($\mathrm{Homeo}(\mathcal{A})$) that reverses the orientation of that circle but not of the other circles of the path (arc); all such elements belong to distinct cosets. And $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ (respectively $\mathrm{Homeo}^{+}(\mathcal{A})$) is not normal in $\mathrm{Homeo}(\mathcal{R}_{n})$ (respectively $\mathrm{Homeo}(\mathcal{A})$, the orientation depends on the embedding of the Julia set in $\mathbf{C}$.
@@ 3.3 Simplicity of commutator subgroups
Figure 9: The 33-coloring of the circles of $\mathcal{R}_{3}$.
@@
The basilica and airplane rearrangement groups $T_{B}$ and $T_{A}$ from [BF15, BF19, Tar24a] are countable dense subgroups of the groups $\mathrm{Homeo}^{+}(\mathcal{B})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$, respectively. For the basilica this was shown in [Ner24] and for the airplane a proof is sketched in [Tar25, Subsection 7.3].
Note that the behavior of $\mathrm{Homeo}^{+}(\mathcal{B})$ described in Theorem 3.9 is shared by $T_{B}$, since its commutator subgroup $T_{B},T_{B}]$ is a simple index-$2$ normal subgroup consisting of those elements of $T_{B}$ that preserve the $2$-coloring depicted in Figure 8 [BF15, Section 8]. Instead, unlike $\mathrm{Homeo}^{+}(\mathcal{A})$, the airplane rearrangement group $T_{A}$ has infinite-index simple commutator subgroup [Tar24a, Corollary 7.4 and Thoerem 7.8].
@@
    A rearrangement group of the basilica that does not preserve the orientation of the circles can be built by switching orientation of the edge attached to the initial vertex of the replacement graph of the basilica replacement system. The same arguments employed in [Ner24, Tar25] show that it is dense in $\mathrm{Homeo}(\mathcal{B})$. Does it have a simple index-$2$ commutator subgroup?
@@ 4.1 Patchwork
Since most of the statements are essentially the same in both settings, we will let $T$ be a biregular tree and $D$ be a Ważewski dendrite. Given $p$ and $q$ vertices of a tree or points of a dendrite, we will denote by $[p,q]$ either the unique path in the tree (as a set of vertices) or the unique arc in the dendrite (as a subset of the dendrite) between $p$ and $q$. Recall that, given $x$, $y$ and $z$ distinct vertices of $T$ or points of $D$, their center is the unique point in the intersection $x,y]\cap[y,z]\cap[z,x]$.
@@
  • Given subsets $F$ and $F^{\prime}$ of $V(T)$, a partial tree homomorphism $F\to F^{\prime}$ is a bijection that is a restriction of a tree isomorphism $F]\to[F^{\prime}]$.
@@
It is easy to see that any partial tree homomorphism $F\to F^{\prime}$ is the restriction of a unique tree isomorphism $F]\to[F^{\prime}]$.
@@
If $F$ is a center-closed subset of a dendrite $D$, then each element of $\Omega_{F}$ has at most two points of $F$ on its topological boundary. Indeed, if $\partial A\supseteq\{x,y,z\}$ for some $A\in\Omega_{F}$, then the center of $x$, $y$ and $z$ would belong to $F$ and so $A$ could not be a connected component of $D\setminus F$.
@@
Suppose that $F$ and $F^{\prime}$ are subsets of $V(T)$ or $D$ and let $f\colon F\to F^{\prime}$ be a partial homomorphism. In the case of a dendrite, suppose also that $F$ and $F^{\prime}$ are closed, center-closed and that, for each two cut points $x$ and $y$ of $D$, $x,y]\cap F$ and $x,y]\cap F^{\prime}$ are finite. Assume the following:
@@
Define a map $h\colon V(T)\to V(T)$ or $D\to D$ by setting $h(x)=h_{A}(x)$ when $x\in A$ for some $A\in\Omega_{F}$ and $h(x)=f(x)$ for all $x\in F$. Then $h$ is an automorphism of $T$ or a homeomorphism of $D$ that agrees with $f$ on $F$.
Note that the additional hypotheses required on $F$ and $F^{\prime}$ in the case of a dendrite always hold when such sets are finite and center-closed.
@@
For dendrites, this is Lemma 4.3 from [DMW19] when $F$ and $F^{\prime}$ are finite. When they are infinite, we will show that $h$ is continuous by proving that it maps connected components of the complement of a point to connected components of the complement of a point, thus preserving a subbasis for the topology of the dendrite. This is equivalent to showing that $h$ maps arcs to arcs. Observe that each $A\in\Omega_{F}$ or $B\in\Omega_{F^{\prime}}$ has non-empty topological boundary. Indeed, $A$ is open because $D$ is locally connected and $F$ is closed, so if $\partial A$ were empty then $A$ would be clopen, which contradicts the fact that $D$ is connected (and the same for any $B$ in $\Omega_{F^{\prime}}$). Then $\partial A$ (respectively $\partial B$) consists exactly of one or two points of $F$ (respectively $F^{\prime}$) by Remark 4.3. Now, each arc $[x,y]$ of $D$ with $x$ and $y$ cut points splits into finitely many closed subarcs with end points in $F\cup\{x,y\}$, so $h$ maps it to an arc, because it is a homeomorphism $h_{A}$ on the interior of each subarc and because two adjacent subarcs meet in a point of $F$ on which $h$ has the same value as $f$. If $x$ is not a cut point but $y$ is (the case when neither are cut points is similar), then $(x,y]$ is an increasing union of arcs $x_{i},y_{i}]$ whose end points are cut points, so the previous argument applies to each $x_{i},y_{i}]$, thus on $(x,y]$, and ultimately to $[x,y]$ because if $x\in F$ then $h(x)=f(x)$ and if $x\notin F$ then $h(x)=h_{A}(x)$ for some $A\in\Omega_{F}$.
Let us now consider the case of a tree. In this case, first note that $h$ is a well-defined bijection because $\Omega_{F}\cup\{F\}$ is a partition of $V(T)$. Then we only need to check that the bijection $h$ of the vertices of $T$ is truly a tree automorphism, i.e., that $\{v,w\}\in E(T)$ if and only if $\{h(v),h(w)\}\in E(T)$. When both $v$ and $w$ belong to $F$, this follows from the fact that $h|_{F}=f|_{F}$ and that $f$ descends from an isomorphism $F]\to[F^{\prime}]$. On the other hand, when either or both vertices belong to the complement of $F$, since they are at distance $1$ they must belong to a unique common $A\cup\partial A$, so the conclusion follows from the fact that each $h_{A}$ is a tree automorphism defined on $A\cup F$. ∎